OpenStudy (clara1223):

simplify 8sec(x)/(tan(x)+cot(x))

OpenStudy (clara1223):

\[\frac{ 8\sec(x) }{ \tan(x)+\cot(x) }\]

OpenStudy (clara1223):

a) 8cot(x) b) 8cos(x) c) 8csc(x) d) 8sec(x) e) 8sin(x)

OpenStudy (clara1223):

I have gotten it to \[8\frac{ \sin(x) }{ \cos ^{2}(x) }\] can anyone at least simplify from there?